Please help me with the folowing....
How do you solve for each variable:
(x+4)(2x-9)=0
x^2-3x-10=0
4x^2-25=0
10=x^2-9x
9c^2=36

Answer by checkley77(12844)

(x+4)(2x-9)=0
x+4=0
x=-4 answer
2x-9=0
2x=9
x=9/2
x=4.5
------------------------------
x^2-3x-10=0
(x-5)(x=2)=0
x-5=0
x=5 answer.
x+2=0
x=-2 answer.
-----------------------------------------
I've done the first 2 now you can finish the homework.